### v1

# Gregory Weitzner

Abstract
People are often reluctant to incorporate information produced by algorithms into their decisions, a phenomenon called “algorithm aversion”. This paper shows how algorithm aversion arises when the choice to follow an algorithm conveys information about a human’s ability. I develop a model in which workers make forecasts of a random outcome based on their own private information and an algorithm’s signal. Low-skill workers receive worse information than the algorithm and hence should always follow the algorithm’s signal, while high-skill workers receive better information than the algorithm and should sometimes override it. However, due to reputational concerns, low-skill workers inefficiently override the algorithm to increase the likelihood they are perceived as high-skill. The model provides a fully rational microfoundation for algorithm aversion that aligns with the broad concern that AI systems will displace many types of workers.
1 Introduction
In recent years, algorithms have become better at forecasting many types of outcomes than humans, raising concerns that artificial intelligence will cause the mass displacement of jobs. [^1] I follow Ludwig and Mullainathan (2021) by using the phrase algorithm interchangably with “artificial intelligence”, “machine learning” or “deep learning”. An incomplete list of papers that show algorithms outperform humans across many domains include: Yeomans et al. (2019), Lai et al. (2021), Mullainathan and Obermeyer (2019), Liu (2022) and Agarwal et al. (2023). However, many argue that these concerns are overstated: instead of being displaced by algorithms, humans will play a critical role by collaborating with and/or supervising algorithms. [^2] E.g., Langlotz (2019), Agrawal, Gans, and Goldfarb (2019), Ludwig and Mullainathan (2021) Acemoglu, Autor, and Johnson (2023). For example, Ludwig and Mullainathan (2021) argue that in many contexts, humans can supplement algorithm’s because they have better context/can better assess rare situations by using intuition. Based on this idea, a rapidly growing literature across computer science, economics and psychology explores how humans work alongside artificial intelligence systems.
However, despite the clear value of algorithms, humans are often reluctant to incorporate algorithmic forecasts into their own decisions, a phenomenon which Dietvorst, Simmons, and Massey (2015) coin “algorithm aversion”. Algorithm aversion has been documented empirically in domains as disparate as health (e.g., Promberger and Baron (2006), Longoni, Bonezzi, and Morewedge (2019), Shaffer et al. (2013) and Agarwal et al. (2023)), finance (e.g., Önkal et al. (2009), Niszczota and Kaszás (2020) and Greig et al. (2023)) and joke recommendations (e.g., Yeomans et al. (2019)). In some cases, even after being provided the algorithm’s forecast, humans’ forecasts perform worse than the algorithm on its own. At first blush, this behavior is surprising. If humans act as rational Bayesians, their forecasts should always be at least as accurate as an algorithm’s. For this reason, the main explanations for algorithm aversion center on behavioral biases and psychological reasons. [^3] See Burton, Stein, and Jensen (2020), Jussupow, Benbasat, and Heinzl (2020) and Mahmud et al. (2022) for review articles discussing the various explanations for algorithm aversion. Given the large potential productivity benefits of AI-human collaboration, it is critical to understand why exactly humans are so reluctant to incorporate valuable information from algorithms into their own decision making. Understanding the reasons for algorithm aversion can also help in design more efficient systems in which humans work alongside algorithms.
In this paper, I argue that a simple, yet overlooked, reason for algorithm aversion is that a human’s choice to follow or override an algorithm’s forecast may convey information about that human’s ability. To illustrate this idea, consider a situation in there is worker forecasting an outcome with the help of an algorithm provided by the firm. After observing the algorithm’s forecast, the worker makes his own forecast and the outcome is realized. The worker has private information about their own skill, where he can either be high-skill, in which case he is better at predicting the outcome than the algorithm, or he can be low skill, in which case he is worse than the algorithm. Hence, it is efficient for the low-skill worker to report the algorithm’s forecast, while the high-skill worker report his own information. This paper shows that the efficient use of the algorithm can never be achieved. Why? If the algorithm is used efficiently, only the high-skill worker ever overrides the algorithm. However, if this were the case, the worker would always override the algorithm to convince the firm they are high-skill. Hence, in equilibrium the low-skill worker must at least sometimes inefficiently override the algorithm. Put in more simple terms, in order for a worker to convince his boss that he is worth keeping around, he needs to sometimes override the algorithm, even if doing so is the wrong decision for the company.
Formally, I analyze a strategic communications, i.e., cheap talk, game in which a worker is tasked with predicted an uncertain outcome. The worker receives a private signal as well as a public signal produced by an algorithm. After receiving both signals, the worker reports a forecast, after which the outcome is realized. Importantly, the worker has private information about his own skill. Specifically, the worker can either be low-skill, in which case his information is always inferior to the algorithm, or he can be high-skill in which case his information is always superior to the algorithm. The firm prefers the worker report an accurate forecast; hence, from the firm’s perspective, it is efficient for the low-skill worker to report the algorithm’s forecast, while the high-skill worker always reports their own signal. After receiving the worker’s forecast and the realization of the outcome, the firm updates their beliefs regarding the worker’s skill. Importantly, the worker faces reputational concerns (e.g., Holmström (1999)) and hence reports a forecast which maximizes the firm’s perception that he is highly skilled.
I first consider a benchmark case with no algorithm. In this benchmark case, there always exists an efficient equilibrium in which the worker honestly reports his signal. Intuitively, the firm assesses the worker entirely on the accuracy of his forecast which incentivizes the worker to always reports his signal truthfully in order for his forecast to be perceived as accurate as possible. In the main model in which the worker has access to the algorithm’s signal, this is no longer the case. Here it is efficient for the low-skill worker to report the algorithm’s forecast, while the high-skill worker should always report his own signal. However, whereas in the benchmark case the forecast on its own provided no information regarding the worker’s skill, here the firm learns both from the accuracy of the worker’s forecast and the forecast itself because it can be compared to the algorithm’s signal. In fact, because of this latter effect the first-best use of information can never be achieved. The reason behind this result is simple: if the high-skill worker is the only type to ever defy the algorithm, defying the algorithm will instantly reveal that the worker is high-skill. Hence, under this set of beliefs, both worker types would always deviate by reporting the opposite signal of the algorithm.
I next show that there exists a unique informative equilibrium in which the high-skill worker always reports his signal and the low-skill worker reports his signal with some positive probability when his signal differs from the algorithm. In this equilibrium, the worker exhibits “algorithm aversion”, i.e., he overrides the algorithm even when his own signal is less accurate than the algorithm. Hence, algorithm aversion causes the overall accuracy of the worker’s forecast to be strictly lower than the first-best. In fact, in some cases the accuracy of the worker’s forecast is worse than the algorithm alone and the worker provides at best zero static benefits to the firm.
I next show that algorithm aversion increases in the uncertainty of the algorithm. Intuitively, as the algorithm becomes more uncertain it becomes less costly to defy it because its signal is less accurate. However, there is also an interesting second effect. Because the algorithm is less accurate, and the high-skill worker always reports his own signal, the high-skill worker’s signal differs from the algorithm more often and hence he defies the algorithm more frequently in equilibrium. This increases the signaling value of defying the algorithm for the low-skill worker. Hence, this indirect effect causes an even larger increase in algorithm aversion. Consistent with this result, Agarwal et al. (2023) show that radiologists’ human-AI diagnostic accuracy goes down when the algorithm’s forecast is more uncertain.
Beyond generating a rational microfoundation for algorithm aversion, the model’s predictions are consistent with substantial empirical evidence. First, humans’ forecasts are often less accurate than an algorithm’s even after being given access to the algorithm’s forecast (e.g., Angelova, Dobbie, and Yang (2023) and Agarwal et al. (2023)), which can occur enogenously in my model for completely rational reputational reasons. Relatedly, Angelova, Dobbie, and Yang (2023) analyze judges’ discretion over algorithm and find that 90% of judges underperform an algorithm when they override it’s recommendation. In my model, the worker overrides the algorithm too frequently and the accuracy of his forecast in the case of an override is often lower than that of the algorithm. While part of Angelova, Dobbie, and Yang (2023)’s results could be due to judges simply making mistakes, my model shows how override decisions can also be driven by reputational concerns. If a worker never override an algorithm, it can be difficult for them to justify keeping their job.
The main mechanism of my model is that by relying on an algorithmic recommendation, a human conveys negative information about their ability. Consistent with this prediction, Shaffer et al. (2013) show that patients rate doctors using computerized recommendations lower than those that do not rely on computerized recommendations. [^4] Leyer and Schneider (2019) show that humans are less likely to delegate decisions to an AI system versus another human. Although there is no delegation decision in my model, if such a delegation decision reflects negatively on the human, my model would also predict a reluctance to do so. Similarly, Jago (2019) show that people evaluate the same piece of art as worse if they learn it was produced by an algorithm. My model would predict that purely learning the the fact that the art was produced by an algorithm can be perceived as a negative signal. [^5] This is assuming there is some “common value” component of the art that is unobservable to the viewer. For example, telling someone a piece of art is made by Van Gogh would in most cases be a positive signal.
To my knowledge this is the first model providing a mechanism for algorithm aversion which is consistent with rational Bayesian behavior. A non-exhaustive list of other explanations include i) the desire to understand forecasts (e.g., Yeomans et al. (2019)), ii) valuing agency of decisions (e.g., Sunstein (2023)), iii) humans having incorrect priors (e.g., Yeomans et al. (2019), Greig et al. (2023)) and iv) humans deriving utility from interacting with another humanan and/or disutility from interacting with a machine (e.g., Dietvorst, Simmons, and Massey (2015) and Greig et al. (2023)). [^6] See Jussupow, Benbasat, and Heinzl (2020), Burton, Stein, and Jensen (2020) and Mahmud et al. (2022) for review articles of algorithm aversion. While I do not doubt that psychological and behavioral reasons may partially explain this phenomenon, I argue that a simple, missing part of the story is the reputational concerns of the people using these algorithms. More broadly, by highlighting an impediment of humans working alongside AI systems, this paper contributes to the literature on human-AI collaboration. [^7] See Lai et al. (2021) for an excellent review of this literature.
In the model, algorithm aversion makes the worker’s forecast less accurate and in some cases the worker provides zero value to the firm in the short-run. In practice, this can put downward pressure on wages for humans. However, if firms are are impatient and it is costly to retain/train workers, reputational algorithm aversion could result in firms firing workers entirely. For these reasons, my paper relates to the literature analyzing the use of AI within firms and in AI’s affect on labor market outcomes (e.g., Acemoglu and Restrepo (2018), Acemoglu and Restrepo (2020), Acemoglu et al. (2022), Cao et al. (2023), Babina et al. (2024), Jha et al. (2024) and Eisfeldt, Schubert, and Zhang (2023)).
Finally, from a theory perspective, this paper builds on the cheap-talk literature first developed by Crawford and Sobel (1982). More specifically, my model can be categorized as reputational cheap-talk (e.g., Scharfstein and Stein (1990), Ottaviani and Sørensen (2006a) Ottaviani and Sørensen (2006b), Ottaviani and Sørensen (2006c), Guembel and Rossetto (2009)) in which agents make strategic forecasts in order to maximize the likelihood that they are perceived as skilled. There are two main differences between my model and typical reputational cheap-talk models. First, in my model the worker knows his type, whereas typically the sender’s type is unknown to himself. Second, the worker has access to both a private signal and a public signal (i.e., the algorithm). The public signal causes the worker’s forecast itself to convey information about his ability beyond the accuracy of the forecast. This causes workers to place too little weight on the public signal due to their reputational concerns.
2 Model Setup
There is a worker who is tasked by his employer to predict an outcome regarding a state of the world $\omega\in\{\omega_{0},\omega_{1}\}$. The worker is of type $\theta\in\{\theta_{L},\theta_{H}\}$ (low-skill or high-skill) and receives a private signal $s\in\{s_{0},s_{1}\}$ regarding $\omega$, where both the worker’s type and signal are only known by the worker. The worker is also given access to an algorithm developed by the firm, which produces a signal $a\in\{a_{0},a_{1}\}$ regarding $\omega$, where the algorithm’s signal is observable to both the worker and the employer. I assume the algorithm is a “black-box” which obtains information regarding $\omega$ by analyzing a set of training data which I do not explicitly model.
The worker observes both signals then reports a message $m\in\{m_{0},m_{1}\}$ to the employer, where $m=m_{0}$ means (in equilibrium) that the worker predicts that $\omega=\omega_{0}$. [^8] Because this is a cheap-talk model, the messages only have meaning to the extent they convey information to the manager in equilibrium.
Because the worker’s actions are costless, the model applies to any situation in which the worker’s action acts as a forecast which conveys information about his skill. For example, it could be the worker is forecasting the firm’s earnings or cash flows, but could also be be that the worker is making an investment decision (e.g., Scharfstein and Stein (1990)). [^9] In Scharfstein and Stein (1990) the outcome of the investment decision is observable whether or not the firm invests. This could be relevant to situations in which the worker is tasked with predicting some type of market-wide conditions.
2.1 Information Structure
The worker is equally likely to be low-skill or high-skill, i.e., $Pr(\theta_{H})=Pr(\theta_{L})=\frac{1}{2}$ and both states of the world are equally likely, $Pr(\omega_{1})=Pr(\omega_{0})=\frac{1}{2}$. When $\omega=\omega_{1}$, the high-skill worker receives the signal $s_{1}$ with probability $\upsilon_{H}$, while the low-skill worker receives it with probability $\upsilon_{L}<\upsilon_{H}$. When $\omega=\omega_{0}$, the high-skill worker receives the signal $s_{1}$ with probability $1-\upsilon_{H}$ and the low-skill worker receives it with probability $1-\upsilon_{L}$.
$Pr(s_{1}|\omega_{1},\theta_{L})$ $\equiv\upsilon_{L}>\frac{1}{2},$ (1)
$Pr(s_{1}|\omega_{1},\theta_{H})$ $\equiv\upsilon_{H}>\upsilon_{L},$ (2)
$Pr(s_{1}|\omega_{0},\theta_{L})$ $\equiv 1-\upsilon_{L},$ (3)
$Pr(s_{1}|\omega_{0},\theta_{H})$ $\equiv 1-\upsilon_{H}.$ (4)
These assumptions make it such that the high-skill worker always has better information than the low-skill worker and that the ex-ante distribution of signals is the same for both low-skill and high-skill workers, i.e., $Pr(s_{1}|\theta_{H})=Pr(s_{1}|\theta_{L})=\frac{1}{2}$. The latter ensures that the signal itself does not convey any information about the skill of the worker. [^10] A similar assumption is made in Scharfstein and Stein (1990). However, as shown below, the employer can obtain information about the worker by comparing his forecast to the algorithm’s signal.
When $\omega=\omega_{1}$, the algorithm produces the signal $a_{1}$ from the algorithm with probability $\alpha$, while if $\omega=\omega_{0}$ the algorithm produces signal $a_{1}$ with probability $1-\alpha$. Importantly, the algorithm’s recommendation does not depend on the worker’s skill. This assumption is reasonable so long as the algorithm is developed by other employees of the firm or outside the firm entirely. In practice, different firms may even use the same algorithm developed by a third party. [^11] For example, firms often us the same large language models, e.g., ChatGPT. Notice that the errors in the worker signals and algorithm signals are conditionally independent. This helps simplify the expressions without meaningfully affecting the main mechanism.
To make the problem interesting, I assume that the high-skill worker’s signal is more informative than the algorithm, while the low skill worker’s is less informative.
Assumption 1.
the high-skill worker’s signal is more informative than the algorithm, while the low skill worker’s is less informative, i.e., $\alpha\in(\upsilon_{L},\upsilon_{H})$.
Without this assumption, it would always be optimal for the worker to either always report the algorithm’s signal or always report their own signal. In practice, the “ground truth” in algorithms’ training data is often established based on “expert” assessments. Similarly, Angelova, Dobbie, and Yang (2023) find that 10% of judges outperform an algorithm in the context of bail decisions. Hence, it is natural that at least some workers can forecast better than the algorithm. Moreover, if the algorithm always provided the better forecast, there would be no need to employ the worker at all.
2.2 Worker’s Objective and Equilibrium Concept
I assume that the manager always prefers that the worker’s forecast is correct but do not explicitly model why this is valuable to the manager. [^12] This is standard in reputational cheap talk models (e.g., Scharfstein and Stein (1990), Ottaviani and Sørensen (2006a), Ottaviani and Sørensen (2006b), Ottaviani and Sørensen (2006c). For example, the firm may prefer that it’s treasurer make more accurate cash flow estimates to help guide firms’ financial policies. Relatedly, the firm prefers to make investments when market conditions turn out to be favorable for that investment.
The manager first observes the worker’s forecast $m$ and the algorithm’s signal $a$, after which the outcome $\omega$ is realized. The manager then forms a posterior belief $\hat{\theta}(m,a,\omega)$ on the probability that the worker is high-skill. [^13] We can equivalently think of the manager as anyone who observes both the worker and the algorithm’s recommendation. If the algorithm is entirely public, then we think of the manager as the entire labor market (e.g., Holmström (1999)).
Through assessing the worker’s forecast and the accuracy of the forecast in predicting the outcome, the manager can learn about the worker’s skill. As is typical in reputational models, there are no long-term contracts and hence, the worker attempts to maximize the manager’s perceived probability that the worker is skilled. This is equivalent to a risk-neutral worker maximizing his next period, competitive spot wage, where the wage is linear in the manager’s posterior belief that the worker is high-skill. [^14] This approach is common in reputational cheap talk models (e.g., Scharfstein and Stein (1990)). Hence, the worker’s payoff following his own signal $s$, the algorithm’s signal $s$ and the worker’s message $m$ as follows:
$\sum_{\omega}Pr(\omega|s,\theta)\hat{\theta}(m,a,\omega).$ (5)
The equilibrium concept I use throughout the analysis is perfect Bayesian Nash Equilibrium, in which the worker’s equilibrium strategy maximizes his expected reputational payoff and the employer’s conditional probabilistic belief is correct on equilibrium path, i.e., $\hat{\theta}(m,a,\omega)=Pr(\theta_{H}|m,a,\omega)$ for $\omega\in\{\omega_{0},\omega_{1}\}$, $a\in\{a_{0},a_{1}\}$ and any $m$ that is used in equilibrium.
3 Benchmark: No Algorithm
First, it will be useful to establish a benchmark in which the worker does not have access to the algorithm’s signal to highlight the mechanism throughwhich algorithm aversion arises. Because the employer always prefers a more accurate forecast, it is efficient for the worker to report his signal truthfully. For this analysis, it will be useful to calculate the worker’s posterior belief regarding the likelihood of state $\omega$:
$Pr(\omega_{1}|s_{1},\theta_{H})=\frac{Pr(s_{1}|\omega_{1},\theta_{H})Pr(\omega_{1}|\theta_{H})}{Pr(s_{1}|\theta_{H})}$ $=\upsilon_{H},$ (6)
$Pr(\omega_{1}|s_{0},\theta_{H})=\frac{Pr(s_{0}|\omega_{1},\theta_{H})Pr(\omega_{1}|\theta_{H})}{Pr(s_{0}|\theta_{H})}$ $=1-\upsilon_{H},$ (7)
$Pr(\omega_{1}|s_{1},\theta_{L})=\frac{Pr(s_{1}|\omega_{1},\theta_{L})Pr(\omega_{1}|\theta_{L})}{Pr(s_{1}|\theta_{L})}$ $=\upsilon_{L},$ (8)
$Pr(\omega_{1}|s_{0},\theta_{L})=\frac{Pr(s_{0}|\omega_{1},\theta_{L})Pr(\omega_{1}|\theta_{L})}{Pr(s_{0}|\theta_{L})}$ $=1-\upsilon_{L}.$ (9)
It will also be useful to calculate the manager’s posterior beliefs assuming that the worker engages in truth-telling:
$\hat{\theta}(\omega_{1},s_{1})$ $=\frac{\upsilon_{H}}{\upsilon_{H}+\upsilon_{L}},$ (10)
$\hat{\theta}(\omega_{1},s_{0})$ $=\frac{1-\upsilon_{H}}{2-\upsilon_{H}-\upsilon_{L}},$ (11)
$\hat{\theta}(\omega_{0},s_{1})$ $=\frac{1-\upsilon_{H}}{2-\upsilon_{H}-\upsilon_{L}},$ (12)
$\hat{\theta}(\omega_{0},s_{0})$ $=\frac{\upsilon_{H}}{\upsilon_{H}+\upsilon_{L}}.$ (13)
In order for the low-skill and high-skill workers to truthfully report their signals, the following incentive compatibility conditions must hold:
$\sum_{\omega}Pr(\omega|s_{1},\theta_{L})\hat{\theta}(s_{1},\omega)$ $\geq\sum_{\omega}Pr(\omega|s_{1},\theta_{L})\hat{\theta}(s_{0},\omega)$ (14)
$\sum_{\omega}Pr(\omega|s_{1},\theta_{H})\hat{\theta}(s_{1},\omega)$ $\geq\sum_{\omega}Pr(\omega|s_{1},\theta_{H})\hat{\theta}(s_{0},\omega).$ (15)
I next show that in this benchmark case, there always exists an equilibrium in which the worker reports his signal as his forecast. {prop} In the benchmark case without the algorithm, a truth-telling equilibrium always exists in which both the low-skill worker and high-skill worker truthfully report their signal, i.e., $m=m_{0}$ if $s=s_{0}$ and $m=m_{1}$ if $s=s_{1}$.
Proof.
Because the problem is symmetric, the incentives to deviate will always be the same regardless of whether the worker receives a good or bad signal. Hence, it is without loss of generality to consider the incentives to deviate following a good signal. Plugging in the posteriors calculated above into (14) and simplifying we have:
$\frac{\left(\upsilon_{H}-\upsilon_{L}\right)\left(2\upsilon_{L}-1\right)}{\left(2-\upsilon_{H}-\upsilon_{L}\right)\left(\upsilon_{H}+\upsilon_{L}\right)},$ (16)
which is positive given that $\upsilon_{H}>\upsilon_{L}$ and $\upsilon_{L}>\frac{1}{2}$. Hence, the low-skill worker will not deviate. Plugging in the posteriors into (14) and simplifying we have:
$\frac{\left(\upsilon_{H}-\upsilon_{L}\right)\left(2\upsilon_{H}-1\right)}{\left(2-\upsilon_{H}-\upsilon_{L}\right)\left(\upsilon_{H}+\upsilon_{L}\right)},$ (17)
[…]
Intuitively, because the worker’s forecast itself does not convey any information about the worker’s ability, the worker can only convey his skill by making as accurate of a forecast as possible. Hence, it is incentive compatible for the worker to report his signal truthfully if his manager expects him too. However, once I introduce the algorithm, the worker’s forecast will convey information about his skill because it can be compared to the algorithm’s signal.
It is worth mentioning that as in all cheap talk models, there also exists a “babbling” equilibrium, in which both high and low-skill workers report a forecast $m$ with the same probability regardless of their type and the signal they receive. In this case, the employer cannot glean any information from the worker’s message and hence, the worker is indifferent between any potential message he sends.
4 Algorithm
Next, I analyze the case in which the worker has access to the algorithm before making his forecast. Because the problem is symmetric, throughout the analysis I focus on the case in which the worker receives signal $s_{1}$; however, the analysis is the same for the case in which $s=s_{0}$. Again, it will be useful to calculate the worker’s posterior belief regarding the likelihood of the state:
$Pr(\omega_{1}|s_{1},a_{1},\theta_{L})=\frac{Pr(s_{1},a_{1}|\omega_{1},\theta_{H})Pr(\omega_{1}|\theta_{H})}{Pr(s_{1},a_{1}|\theta_{H})}$ $=\frac{\alpha\upsilon_{L}}{\alpha\upsilon_{L}+(1-\alpha)(1-\upsilon_{L})},$ (18)
$Pr(\omega_{1}|s_{1},a_{1},\theta_{H})=\frac{Pr(s_{1},a_{1}|\omega_{1},\theta_{H})Pr(\omega_{1}|\theta_{H})}{Pr(s_{1},a_{1}|\theta_{H})}$ $=\frac{\alpha\upsilon_{H}}{\alpha\upsilon_{H}+(1-\alpha)(1-\upsilon_{H})},$ (19)
$Pr(\omega_{1}|s_{1},a_{0},\theta_{L})=\frac{Pr(s_{1},a_{0}|\omega_{1},\theta_{L})Pr(\omega_{1}|\theta_{L})}{Pr(s_{1},a_{0}|\theta_{L})}$ $=\frac{(1-\alpha)\upsilon_{L}}{(1-\alpha)\upsilon_{L}+\alpha(1-\upsilon_{L})},$ (20)
$Pr(\omega_{1}|s_{1},a_{0},\theta_{H})=\frac{Pr(s_{1},a_{0}|\omega_{1},\theta_{H})Pr(\omega_{1}|\theta_{H})}{Pr(s_{1},a_{0}|\theta_{H})}$ $=\frac{(1-\alpha)\upsilon_{H}}{(1-\alpha)\upsilon_{H}+\alpha(1-\upsilon_{H})}.$ (21)
Based on these posterior probabilities, the manager would always prefer that the low-skill worker report the algorithm’s forecast rather than his signal, while the high-skill worker reports his own signal. To see this, first note that from Assumption 1, $\upsilon_{H}>\alpha>\upsilon_{L}>\frac{1}{2}$. Hence, when the worker’s signal coincides with that of the algorithm, the posterior probability that the outcome coincides with the worker’s signal is always greater than one half, i.e., $Pr(\omega_{1}|s_{1},a_{1},\theta)>\frac{1}{2}$ for $\theta\in\{\theta_{L},\theta_{H}\}$. However, when the worker’s signal differs from the algorithm, the high-skill worker’s posterior probability that the outcome coincides with the signal is greater than one-half, i.e., $Pr(\omega_{1}|s_{1},a_{0},\theta_{H})>\frac{1}{2}$, while the low skill’s worker is less than one-half, i.e., $Pr(\omega_{1}|s_{1},a_{0},\theta_{L})<\frac{1}{2}$. Hence, the manager always prefers that the high-skill worker report his own signal, while the low-skill worker always report the algorithm’s signal.
Henceforth, I refer to the case in which the worker reports the algorithm’s signal when he is low-skill and his own signal when he is high-skill the “first-best”. I next show that because of reputational concerns, the first-best can never be achieved as an equilibrium. {prop} The first-best cannot be achieved as an equilibrium.
Proof.
First, it is useful to calculate the following posterior beliefs given that the manager expects the worker to follow the first-best use of information:
$\hat{\theta}^{FB}(m_{1},a_{1},\omega_{1})$ $=\frac{\upsilon_{H}}{\upsilon_{H}+1},$ (22)
$\hat{\theta}^{FB}(m_{1},a_{1},\omega_{0})$ $=\frac{1-\upsilon_{H}}{1-\upsilon_{H}+1},$ (23)
$\hat{\theta}^{FB}(m_{0},a_{1},\omega_{1})$ $=1,$ (24)
$\hat{\theta}^{FB}(m_{1},a_{1},\omega)$ $=1.$ (25)
After receiving a signal $s_{1}$, incentive compatibility for the low-skill worker requires:
$\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{L})\hat{\theta}^{FB}(m_{1},a_{1},\omega)$ $\geq\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{L})\hat{\theta}^{FB}(m_{0},a_{1},\omega),$ (26)
$\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L})\hat{\theta}^{FB}(m_{0},a_{0},\omega)$ $\geq\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L})\hat{\theta}^{FB}(m_{1},a_{0},\omega),$ (27)
where (26) says that when the low-skill worker receives the signal $s_{1}$ he should report $m_{1}$ when the algorithm’s signal is $a_{1}$ and $m_{0}$ when the algorithm’s signal is $a_{0}$. Plugging in the first-best posterior beliefs we have:
$\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{L})\frac{\upsilon_{H}}{\upsilon_{H}+1}$ $\geq\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{L}),$ (28)
$\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L})\frac{1-\upsilon_{H}}{1-\upsilon_{H}+1}$ $\geq\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L}).$ (29)
[…]
Intuitively, under the first-best only the high-skill worker ever overrides the algorithm. Because of this, by overriding the algorithm any worker will be perceived as high-skill with probability one. Hence, both the low-skill and high-skill workers would have incentives to override the algorithm no matter what their signal is in order to be perceived as high-skill. As shown below, there exists an equilibrium in which the low-skill worker sometimes overrides the algorithm even though it is inefficient for him to do so. Before presenting the proposition, I establish the following definition regarding an “informative equilibrium”, in which some information is conveyed regardless of the algorithm’s outcome:
Definition 1.
In an informative equilibrium, the low and high-skill workers do not report the same message with probability one following either algorithmic signal, i.e., when $a=a_{0}$ and when $a=a_{1}$.
Based on this definition, I next prove the following lemma which says the high-skill worker will always report his signal in any informative equilibrium.
Lemma 1.
In any informative equilibrium, the high-skill worker always reports his signal.
Proof.
Suppose to the contrary the high-skill worker always reports the the opposite of his own signal. Again focusing on the case in which the worker receives signal $s_{1}$, first consider the case when the algorithm’s signal is $a_{1}$. For the high-skill worker to report $m_{0}$, the following condition must hold:
$\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{H})\hat{\theta}(m_{1},a_{1},\omega)$ $\leq\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{H})\hat{\theta}(m_{0},a_{1},\omega),$ (30)
Which we can rewrite as:
$Pr(\omega_{1}|s_{1},a_{1},\theta_{H})\left(\hat{\theta}(m_{1},a_{1},\omega_{1})-\hat{\theta}(m_{0},a_{1},\omega_{1})\right)$ (31)
$\leq Pr(\omega_{0}|s_{1},a_{1},\theta_{H})\left(\hat{\theta}(m_{0},a_{1},\omega_{0})-\hat{\theta}(m_{1},a_{1},\omega_{0})\right).$ (32)
However, if (31) holds, the low-skill worker would also always report $m_{0}$ when $s=s_{1}$ and $a=a_{1}$ because $Pr(\omega_{1}|s_{1},a_{1},\theta_{H})>Pr(\omega_{1}|s_{1},a_{1},\theta_{L})$ and $Pr(\omega_{0}|s_{1},a_{1},\theta_{L})>Pr(\omega_{0}|s_{1},a_{1},\theta_{H})$. Hence, both the low and high-skill workers send the message $m_{0}$ when $a=a_{1}$, which is a contradiction that the equilibrium is informative.
Now consider the case in which the algorithm’s signal is $a_{0}$. For the high-skill worker to report $m_{0}$ when his signal is $s_{1}$ we have the following condition:
$Pr(\omega_{1}|s_{1},a_{0},\theta_{H})\left(\hat{\theta}(m_{1},a_{0},\omega_{1})-\hat{\theta}(m_{0},a_{0},\omega_{1})\right)$ (33)
$<Pr(\omega_{0}|s_{1},a_{0},\theta_{H})\left(\hat{\theta}(m_{0},a_{0},\omega_{0})-\hat{\theta}(m_{1},a_{0},\omega_{0})\right).$ (34)
[…]
Following Lemma 1, the following proposition characterizes the unique informative equilibrium.
{prop}
[Algorithm Aversion] There exists a unique, informative equilibrium in which the high-skill worker always reports his own signal and the low-skill worker reports the algorithm’s signal with probability $\gamma\in(0,1)$ and his own signal with probability $1-\gamma$, where $\gamma$ is the unique solution to to the following equation:
$\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L})\hat{\theta}(m_{0},a_{0},\omega)-\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L})\hat{\theta}(m_{1},a_{0},\omega)=0,$ (35)
where
$\hat{\theta}(m_{1},a_{0},\omega_{1})$ $=\frac{\upsilon_{H}}{\upsilon_{H}+(1-\gamma)\upsilon_{L}},$ (36)
$\hat{\theta}(m_{1},a_{0},\omega_{0})$ $=\frac{1-\upsilon_{H}}{1-\upsilon_{H}+(1-\gamma)\left(1-\upsilon_{L}\right)},$ (37)
$\hat{\theta}(m_{0},a_{0},\omega_{1})$ $=\frac{1-\upsilon_{H}}{\left(1-\upsilon_{H}\right)+\left(1-\upsilon_{L}\right)+\gamma\upsilon_{L}},$ (38)
$\hat{\theta}(m_{0},a_{0},\omega_{0})$ $=\frac{\upsilon_{H}}{\upsilon_{H}+\upsilon_{L}+\gamma\left(1-\upsilon_{L}\right)}.$ (39)
Proof.
Because of symmetry, we can again restrict focus to the case when the worker receives signal $s_{1}$. From Lemma 1 we know that in any informative equilibrium the high-skill worker reports the same message as his signal. First, consider the case in which the low-skill worker receives the same signal as the algorithm, i.e., $a=a_{1}$. Here, we need to check whether the low-skill worker has incentives to report $m_{0}$:
$\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{L})\hat{\theta}(m_{1},a_{1},\omega)\geq\sum_{\omega}Pr(\omega|s_{1},a_{1},\theta_{L})\hat{\theta}(m_{0},a_{1},\omega).$ (40)
From Lemma 1 the high-skill worker would never deviate in an informative equilibrium, which again implies that (40) holds. Now consider, the case in which $a=a_{0}$. The low-skill worker can report his own signal or the algorithm’s with some probability $\gamma$. We can calculate the manager’s posterior beliefs as follows:
$\hat{\theta}(m_{1},a_{0},\omega_{1})$ $=\frac{\upsilon_{H}}{\upsilon_{H}+\upsilon_{L}+\gamma\left(1-\upsilon_{L}\right)},$ (41)
$\hat{\theta}(m_{1},a_{0},\omega_{0})$ $=\frac{1-\upsilon_{H}}{1-\upsilon_{H}+1-\upsilon_{L}+\gamma\upsilon_{L}},$ (42)
$\hat{\theta}(m_{0},a_{0},\omega_{1})$ $=\frac{1-\upsilon_{H}}{1-\upsilon_{H}+(1-\gamma)\left(1-\upsilon_{L}\right)},$ (43)
$\hat{\theta}(m_{0},a_{0},\omega_{0})$ $=\frac{\upsilon_{H}}{\upsilon_{H}+(1-\gamma)\upsilon_{L}},$ (44)
where the posteriors do not depend on $a$ because the algorithm’s signal is independent of the worker’s type $\theta$ and independent of the worker’s signal $s$ conditional on the state $\omega$. I define $G(\gamma)$ as the low-skill worker’s difference in payoffs from reporting $m_{1}$ versus $m_{0}$ when he receives signal $s_{1}$ and the algorithm’s signal is $a_{0}$:
$G(\gamma)\equiv\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L})\hat{\theta}(m_{0},a_{0},\omega)-\sum_{\omega}Pr(\omega|s_{1},a_{0},\theta_{L})\hat{\theta}(m_{1},a_{0},\omega).$ (45)
First note from Proposition 4, when $\gamma=1$, we are in the first-best use of information and $G(\gamma)$ is negative, implying that both the low-skill and high-skill worker will always report $m_{1}$. Hence $\gamma$ must be less than $1$ in any informative equilibrium. Differentiating $G(\gamma)$ w.r.t to $\gamma$ we have:
$G^{\prime}(\gamma)=-\frac{\frac{(1-\alpha)\upsilon_{H}\upsilon_{L}^{2}}{\left(\upsilon_{H}+(1-\gamma)\upsilon_{L}\right){}^{2}}+\frac{\alpha\left(1-\upsilon_{H}\right)\left(1-\upsilon_{L}\right){}^{2}}{\left(2-\gamma-\upsilon_{H}-(1-\gamma)\upsilon_{L}\right){}^{2}}+\frac{\alpha\upsilon_{H}\left(1-\upsilon_{L}\right){}^{2}}{\left(\gamma+\upsilon_{H}+(1-\gamma)\upsilon_{L}\right){}^{2}}+\frac{(1-\alpha)\left(1-\upsilon_{H}\right)\upsilon_{L}^{2}}{\left(2-\upsilon_{H}-(1-\gamma)\upsilon_{L}\right){}^{2}}}{\alpha-(2\alpha-1)\upsilon_{L}},$ (46)
which is negative, implying that $G(\gamma)$ is decreasing in $\gamma$. When $\gamma=0$ we have:
$G(0)=\frac{\left(\alpha-\upsilon_{L}\right)\left(\upsilon_{H}-\upsilon_{L}\right)}{\left(2-\upsilon_{H}-\upsilon_{L}\right)\left(\upsilon_{H}+\upsilon_{L}\right)\left(\alpha+(2\alpha-1)\upsilon_{L}\right)},$ (47)
[…]
Proposition 1 shows that, despite the algorithm providing a more informative signal than the low-skill worker’s, the low-skill worker sometimes overrides it when their own signal is different than that of the algorithm’s. Intuitively, low-skill workers cannot always follow the algorithm, otherwise overriding the algorithm signals with probability one that the worker is high-skill. In other words, the worker must override the algorithm occasionally in order to convince their manager they are not dispensable. Hence, Proposition 1 shows how completely rational reputational concerns can endogenously create algorithm aversion.
Agarwal et al. (2023) show that radiologists are less likely to follow an algorithm when they believe that the algorithm’s forecast is more uncertain (e.g., Agarwal et al. (2023)). Motivated by this empirical finding, I next show that the more uncertain the algorithm’s signal is, the more likely low-skill worker overrides the algorithm in the unique informative equilibrium. {prop} Algorithm aversion increases in the uncertainty of the algorithm, i.e., $\gamma$ is increasing in $\alpha$ in the informative equilibrium.
Proof.
Applying the implicit function theorem we have:
$\frac{d\gamma}{d\alpha}=\frac{\partial G(\gamma)/\partial\alpha}{\partial G(\gamma)/\partial\gamma},$ (48)
where the denominator is negative from (46) and the numerator is equal to:
$\sum_{\omega}\frac{\partial Pr(\omega|s_{1},a_{0},\theta_{L})}{\partial\alpha}\hat{\theta}(m_{0},a_{0},\omega)-\sum_{\omega}\frac{\partial Pr(\omega|s_{1},a_{0},\theta_{L})}{\partial\alpha}\hat{\theta}(m_{1},a_{0},\omega)$ (49)
$=\frac{\left(1-\upsilon_{L}\right)\upsilon_{L}}{\left(\alpha-(2\alpha-1)\upsilon_{L}\right){}^{2}}\left(\hat{\theta}(m_{0},a_{0},\omega_{0})-\hat{\theta}(m_{0},a_{0},\omega_{1})+\hat{\theta}(m_{1},a_{0},\omega_{1})-\hat{\theta}(m_{1},a_{0},\omega_{0})\right),$ (50)
which is positive because $\hat{\theta}(m_{0},a_{0},\omega_{0})$ is always larger than $\hat{\theta}(m_{0},a_{0},\omega_{1})$ and $\hat{\theta}(m_{1},a_{0},\omega_{1})$ is always larger than $\hat{\theta}(m_{1},a_{0},\omega_{0})$ because the reputational payoff from the correct forecast is always higher than that of the incorrect forecast. Hence, $\frac{d\gamma}{d\alpha}$ is negative and $\gamma$ is increasing in $\alpha$ in the informative equilibrium.
[…]
This result may at first glance seem obvious. If the algorithm is less accurate, the low-skill worker should be less likely to follow it. However, from the manager’s perspective so long as Assumption 1 is still satisfied, it is always efficient for the low-skill worker to report the algorithm’s signal. Moreover, there are two effects here. On the one hand, it is less costly to override the algorithm given that the algorithm’s signal is uninformative. On the other hand, the high-skill worker also becomes more likely to override the algorithm making it more attractive for the low-skill worker to override the algorithm in order to convey he is high-skill. To see this, we can calculate the probability the worker is high-skill if he receives a signal different than that of the algorithm:
$Pr(\theta_{H}|s_{1},a_{0})$ $=Pr(\theta_{H}|s_{1},a_{0},\omega_{1})Pr(\omega_{1}|s_{1},a_{0})+Pr(\theta_{H}|s_{1},a_{0},\omega_{0})Pr(\omega_{0}|s_{1},a_{0})$ (51)
$=\frac{1}{2}\left(\alpha-(2\alpha-1)\upsilon_{H}\right).$ (52)
Differentiating the RHS of (52) with respect to $\alpha$ we have:
$\frac{1}{2}-\upsilon_{H},$ (53)
which is negative. Hence, $Pr(\theta_{H}|s_{1},a_{0})$ is decreasing in $\alpha$, which implies that the less accurate the algorithm’s signal is, the more likely the worker is skilled if his signal differs from the algorithm. This essentially increases the signaling content of defying the algorithm, increasing the frequency the worker defies the algorithm in equilibrium.
Several studies find that even when being able to incorporate an algorithm’s forecast, humans’ forecasts often perform worse than the algorithm on its own (e.g., Agarwal et al. (2023) and Angelova, Dobbie, and Yang (2023)). Next I show that the this can also occur in my model. {prop} When the average accuracy of the worker’s signal is less than that of the algorithm’s signal, i.e., $\frac{1}{2}(\upsilon_{L}+\upsilon_{H})>\alpha$, the expected accuracy of the worker’s forecast can be lower than that of the algorithm.
Proof.
I prove this by example. Again, because of symmetry we can restrict focus to the case in which the worker reports $m_{1}$ and examine how likely $\omega=\omega_{1}$ as compared to when the algorithm produces signal $a_{1}$. [^16] Because of symmetry $m_{1}$ and $a_{1}$ are always realized half the time. For the worker to be more accurate than the algorithm, the following conditio nmust hold:
$Pr(\omega_{1}|m_{1})$ $\geq Pr(\omega_{1}|a_{1}),$ (54)
$\implies Pr(\omega_{1}|m_{1})$ $\geq\alpha.$ (55)
We can rewrite the LHS of (54) as follows:
$=\frac{Pr(m_{1}|\omega_{1})Pr(\omega)}{Pr(m_{1})}$ (56)
$=Pr(m_{1}|\omega_{1},a_{0})Pr(a_{0}|\omega_{1})+Pr(m_{1}|\omega_{1},a_{1})Pr(a_{1}|\omega_{1})$ (57)
$=\frac{1}{2}\left(\alpha\gamma+\upsilon_{H}+(1-\gamma)\upsilon_{L}\right).$ (58)
First, from this expression notice that if the average accuracy of the worker is higher than that of the algorithm, i.e., $\frac{1}{2}(\upsilon_{L}+\upsilon_{H})>\alpha$, then the worker’s forecast is always more accurate than the algorithm. However, when $\frac{1}{2}(\upsilon_{L}+\upsilon_{H})<\alpha$ this may no longer be the case. For example suppose that $\upsilon_{L}=0.55$, $\alpha=0.60$ and $\upsilon_{H}=0.62$. Solving the informative equilibrium numerically, we have that $\gamma=0.0148$. Plugging these values into the expression for the worker’s accuracy we have $0.58537$ which is clearly less than the algorithm alone $(0.60)$. Hence, there are cases in which the algorithm performs better than the worker’s forecast even though the worker has access to the algorithm’s signal.
[…]
5 Model Discussion and Applications
In this section I discuss some of the key assumptions in the model as well as applications of the model.
First, one may wonder whether the problem can be resolved if the manager simply asks the worker to report their own signal then the manager makes his own forecast. However, this ultimately collapses to the same problem. The low-skill worker would still have incentives to at least sometimes report the algorithm’s signal rather than his own. Relatedly, in some contexts it may not be possible for the worker to convey his signal to the manager, but instead must make a forecast or take an action that indirectly conveys that information. [^17] This could occur if the information cannot be substantiated or if the worker is making a decision, such as an investment, that must be implemented by himself, not the manager.
One potential solution is to make the algorithm’s signal only observable to the worker. In this case, the worker’s forecast cannot be compared to the algorithm’s and hence the worker’s forecast cannot convey information to the manager on its own. There are two practical problems with this. First, the algorithm is likely developed by other people in the firm with expertise in artificial intelligence. These people will inevitably want to see the information produced by the algorithm in order to evaluate and continue developing it. Moreover, even if the firm could make the algorithm’s signal only observable to the worker, it may not be optimal for the firm because it prevents them from learning about the quality of that worker.
The manager could also consider hiding or delaying the availability of the algorithm’s signal to the worker until after the worker reports his forecast. Here, the worker’s problem would collapse to the benchmark case without the algorithm. However, it can easily be shown that the accuracy of the forecast in the benchmark case is strictly lower than that of the informative equilibrium. Moreover, even if the manager ultimately made the final decision, he manager cannot distinguish the quality of the worker’s information given that both low and high-skill workers will always report their information. Hence, the manager either will always follow the worker or the algorithm, depending on which has a higher average accuracy. In contrast, when the worker observes the algorithm, the manager can always do at least as well as the algorithm and sometimes better depending on the average accuracy of the worker’s forecast.
In the informative equilibrium the low-skill worker plays a mixed strategy in which he is indifferent between reporting the algorithm’s signal or his own. While this may not seem realistic at first, mixed strategies can also be interpreted as pure strategies with random disturbances (e.g., Harsanyi (1973)). Equivalently, I could assume that the low-skill worker’s signal precision is a continuous random variable with mean $\upsilon_{L}$ in which the worker reports his own signal whenever the signal precision is above some threshold. This simple alteration would lead to essentially the same equilibrium in pure strategies.
In terms of applications, it is important that the decisions/forecasts the worker makes are consequential enough and not too frequent. Otherwise, the firm can learn quickly the worker’s quality and the inefficiency will be small. For example, a natural application is a employee in a firm’s treasury department projecting the future cash flows of the firm or making an investment decision.
It is also important to emphasize that the model applies to situations in which following an algorithm has a reputational cost for a human. In practice, there are some instances in which a human is presented with a forecast either produced by a human or an algorithm and the human is reluctant to follow the algorithm even though there is no obvious reputational cost of doing so (e.g., see Greig et al. (2023) in the case of robo-advising). The model best applies to situations in which an expert has to make some type of forecast or decision based on his own information and an algorithm’s and he is ultimately evaluated by others. For example, the model applies to any situation in which an employee of a firm must incorporate information from an algorithm into his decisions (e.g., cash flow forecasting, investment decisions and financial policies) and those decisions are ultimately evaluated by his manager or the firm as a whole.
6 Conclusion
Rapid advances in artificial intelligence have raised many questions about the future role of human workers in society. While for certain tasks, algorithms may displace humans entirely, for many others tasks, humans will have to work alongside them. This paper shows how humans’ concerns for their own reputations can impede human-AI collaboration. Completely rational humans may engage in “algorithm aversion” and override the algorithm even if it is inefficient to do so. The intuition for this idea is straightforward: humans have to sometimes override algorithms in order to prove they are not replaceable.
Practically, this can lead to one of two problems: either firms become slower to adopt algorithms, or firms forego using humans for a wider variety of tasks than is would be necessary in the absence of humans’ reputational concerns.
The solution to reputational algorithm aversion is not obvious. As mentioned earlier, reputational concerns can be alleviated if the decision makers are also those put in charge of developing the algorithm (e.g., Dietvorst, Simmons, and Massey (2018)). However, in many cases, these decision makers do not have expertise in AI systems. For example, most radiologists or treasurers of firms likely have little understanding of how a deep learning algorithm works. Similarly, only allowing the decision maker to see the algorithm’s signal would also mitigate reputational concerns. However, this would obviously be difficult if the algorithm is developed by other employees of the firm. Moreover, firms may be tempted to compare their workers’ forecasts to the algorithm in order to learn about their workers’ quality.
Taken together, this paper identifies an entirely rational, yet realistic force that can cause humans to be averse to incorporating information generated by algorithms into their own decisions. Moreover, as AI systems become more and more integrated into firms, I argue that analyzing human-AI interactions through the lens of strategic communication games will be extremely fruitful method to understand these interactions.